# UFEval: Unified Fine-grained Evaluation with Task and Aspect Generalization

### v3

@@ Introduction
To assess UFEval’s generalizability stemming from inter-aspect correlations, we conduct evaluations on previously unseen tasks and aspects, termed Out-of-Domain Evaluation. We then evaluate UFEval as LLM-as-a-Judge using public meta-benchmarks. Finally, we explore the effectiveness of Multi-aspect and Multi-task Assessment Learning for evaluators. In the following sections, we present our experimental setup and corresponding results (In-domain and other experimental results are in Appendix D).
@@
Figure 3: Comparison of baselines on the FRA-OOD and FRA-OOD-H. Black sectors show unseen TAs (for aspect generalization), while the rest show seen UAs for unseen tasks (for task generalization).
Method Tasks Generalization Evaluation Aspects Generalization Evaluation
@@
Table 2: Average accuracy on FRA-OOD and FRA-OOD-H. Seen UAs for Unseen Tasks are used to assess task generalization, while Unseen TAs for Unseen Tasks evaluate aspect generalization.
Experimental Setup
Benchmarks.
(1) Out-of-Domain Evaluation: we use self-constructed FRA-OOD and FRA-OOD-H. (2) Evaluation as LLM-as-a-Judge: we select public meta-benchmarks from three types of tasks. For NLG, we use MT-Bench (Zheng et al. 2023), SummEval (Fabbri et al. 2021), and MANS (Guan et al. 2021). Since SummEval and MANS only provide scores, we generate pairwise samples for evaluation through sample pairing. For IU, we use WildVision (Lu et al. 2024), MLLM-as-a-Judge (Chen et al. 2024), and VLRewardBench (Li et al. 2024c). For IG, GenAI-Bench (Li et al. 2024a) Winoground (Thrush et al. 2022), and Pickapic (Kirstain et al. 2023) are selected. All meta-benchmarks contain human annotation. Moreover, we carefully check that the training set contains no overlapping samples with the meta-benchmarks to ensure fair comparison. Detailed descriptions of each meta-benchmark can be found in Appendix C. (3) Multi-Aspect and Multi-Task Assessment Learning: we use FRA-ID to investigate the multi-aspect learning and the meta-benchmarks to verify multi-task synergy.
@@
Table 3: Evaluation as LLM-as-a-Judge for NLG task. “tau” indicates that accuracy is calculated with ties, and “diff” excludes tied pairs when calculating accuracy. Bold and underline indicate the first and second best results.
@@
Table 4: Evaluation as LLM-as-a-Judge for IU tasks. We evaluate baselines across three meta-benchmarks.
Baselines.
We experiment with representative models from two categories: (1) Prompting Models: GPT-4o, Claude-3.5, Qwen2-VL-72B-Instruct (Qwen2VL-72B), and Qwen2-VL-7B-Instruct (Qwen2VL-7B), implemented using the same prompts as UFEval. (2) Fine-tuned Evaluators: Themis-8B (Hu et al. 2024b) (Themis), Auto-J (Li et al. 2023b), Prometheus 2 (Kim et al. 2024a), LLaVA-Critic-7B (Xiong et al. 2024) (LLaVA-Critic), ImageReward (Xu et al. 2023a), and VisionReward (Xu et al. 2024). Notably, ImageReward and VisionReward use different evaluation paradigms from ours. Auto-J cannot assess specific aspects, and Prometheus 2 requires reference answers. Therefore, we use these methods only as baselines in the LLM-as-a-Judge evaluation.
Out-of-Domain Evaluation
Tasks Generalization Evaluation. For task generalization, we employ the seen UAs and unseen tasks from UFEval during testing to control for test variables. The results are shown in the regions outside the black sectors in Figure 3  [^2] We use aspect abbreviations here; the full names can be found in Appendix A.. Detailed scores for each aspect are reported in Appendix D. UFEval demonstrates strong alignment with both GPT-4o and human annotators on all tasks. In the “Tasks Generalization Evaluation” column of Table 2, UFEval achieves overall average accuracies of 85% and 83% on FRA-OOD and FRA-OOD-H, respectively, and outperforms both Themis and LLaVA-Critic on all tasks. Supporting the effectiveness of aspect-level evaluation for cross-task generalization. We also provide several good and bad cases in Appendix G.
Aspects Generalization Evaluation. We use the unseen aspects of unseen tasks in FRA-OOD and FRA-OOD-H to evaluate UFEval’s generalization to aspects. To support the validity of aspect generalization evaluation, we provide ROUGE_L (Wang et al. 2022) tests on unseen aspects in Appendix B. The results are shown in the black sectors of Figure 3. In terms of overall coverage, UFEval outperforms both Themis and LLaVA-Critic, the state-of-the-art evaluators in their respective domains. In the right column of Table 2, IUFEval continues to demonstrate strong performance on the NLG, IU, IG and ITIG, with high consistency with both GPT-4o and human annotators, achieving overall accuracies of 86.2% and 83.2%, respectively. These results demonstrate that even without exposure to unseen aspects during training, UFEval remains effective in assessment. This is largely due to the inherently interrelated nature of aspects, which enables the transfer of similar semantics and evaluation standards.
Evaluation as LLM-as-a-Judge
Method GenAI-Bench Winoground Pickapic
@@
Table 5: Evaluation as LLM-as-a-Judge for IG task. We evaluate baselines across three meta-benchmarks.
Figure 4: Multi-aspect assessment learning results. We train models using different training set combinations and test them on three visual tasks in FRA-ID to validate the effectiveness of multi-aspect learning.
For NLG tasks. The experimental results, shown in Table 3, demonstrate the effectiveness of UFEval in NLG tasks. Specifically, in SummEval and MT-Bench, UFEval outperforms most baselines, with Claude being the sole exception. In the MANA benchmark, UFEval achieves the best performance, with an accuracy of 69.3%. These results highlight UFEval’s strong capability for text understanding evaluation compared to domain-specific models in this field.
For IU tasks. The results are presented in Table 4. We also evaluate model-level consistency on WildVision using Elo ratings and Kendall’s Tau ($\tau$) to compare model rankings (Xiong et al. 2024). Results demonstrate that UFEval outperforms LLaVA-Critic across all benchmarks. LLaVA-Critic focuses on respective datasets training, whereas UFEval uses joint learning across multiple tasks, which leads to relatively superior performance. More comprehensive MLLM-as-a-Judge results are provided in Appendix D.
For IG Evaluations. The experimental results are shown in Table 5. Our evaluator outperforms ImageReward in all benchmarks, achieving accuracies of 65.5%, 65.7%, and 57.3%, respectively. Even compared to the state-of-the-art image evaluator VisionReward, the performance gap is minimal—only 0.9% lower on GenAI-Bench and 0.3% lower on Winogrounded. These results demonstrate the promising capability of UFEval in the image generation task.
Multi-aspect and Multi-task Assessment Learning
@@
Conclusion
This paper proposes UFEval, the first unified fine-grained evaluator with task and aspect generalization. Specifically, we start by building a comprehensive aspect tree that leads to the creation of FRABench, a large-scale, multi-modal, and aspect-level evaluation dataset. FRABench spans four tasks, 28 sub-tasks, and 112 diverse aspects. We then fine-tune an LMM on FRABench to develop UFEval. Our experimental results demonstrate that joint learning across diverse visual tasks and aspects yields significant mutual benefits and generalization capabilities. In future work, we aim to leverage UFEval to automatically construct high-quality preference pair datasets from model outputs for DPO training.
Appendix A Appendix A: Aspect Sources and Definitions
We provide the sources and definitions of aspects under two categories: UAs and TAs, as shown in Tables 10 to 23. For UAs, the “Target” column indicates the applicable output modality: “T’ for Text, “I” for Image, and “ITI” for Interleaved Text-with-Image. For TAs, the “Sub-Task” column specifies the corresponding task type. Additionally, underscores are used to benchmark extended TAs for ITIG. Abbreviations for all aspects are provided in the table.
Appendix B Appendix B: Analysis of the FRABench
@@
More Detailed Information about the Dataset
@@
We provide the information about sub-tasks for NLG and IU in Table 6, as this information is not shown in Figure 2 of the main text. For the tasks in the “Generated” column in Figure 2 of the main text, due to the absence of pairwise data, we use four LMMs with varying performance levels to generate pairwise responses for each query, as illustrated in Table 7.
Aspect Diversity Analysis
Following prior work (Wang et al. 2022; Honovich et al. 2022), we analyze the ROUGE-L distribution between aspects in the training set and the FRA-OOD, respectively. Specifically, we sample pairs of aspects from the training set and compute their ROUGE-L scores. The overall distribution is shown in Figure 9, which indicates that the selected aspects are distinct from one another, confirming the inclusion of diverse and novel aspects in the training set. Additionally, we measure the ROUGE-L scores between aspects by sampling one aspect from the training set and another from the FRA-OOD. As shown in Figure 10, aspects in the training set and those in the FRA-OOD do not overlap, demonstrating the validity of our out-of-domain evaluations.
Position Bias Analysis
Figure 5: Statistics of pairwise data in the training set.
Figure 6: Statistics of pairwise data in the FRAUAs-OOD.
Figure 7: Statistics of pairwise data in the FRA-ID.
Figure 8: Statistics of pairwise data in the FRA-OOD.
Aspect Selection Across Different Sub-tasks
We provide detailed information on the aspects selected for the training set, FRA-ID, FRAUAs-OOD, and FRA-OOD, as presented in Table 9. Additionally, we include statistics regarding the number of evaluation labels used for training and testing in each sub-task. Blue-marked aspects are human-annotated, while the rest are annotated by GPT-4o. The human-annotated data in IG is sourced from the ImageRewardDB dataset (Xu et al. 2023a).
Appendix C Appendix C: Experiment Detail
Fine-tuning Details of UFEval
Detailed Information of Public Meta-benchmarks
We provide detailed descriptions and experimental information for the public meta-benchmarks:
  • SummEval (Fabbri et al. 2021): SummEval evaluates LLMs on text summarization across four aspects: Fluency, Consistency, Coherence, and Relevance. The dataset contains 100 instructions per aspect, each with responses from 16 different LLMs and corresponding human ratings (totaling 6.4k annotated samples). For testing, we randomly generate 9 comparison pairs from the 16 responses per instruction, creating 3.6k evaluation samples.
  • MANS (Guan et al. 2021): MANS evaluates LLMs on the story generation task. Each instruction includes 5 different responses with human annotation, totaling 400 samples. For each instruction, we randomly select two responses to form pairwise data, resulting in 400 evaluation samples.
  • MT-Bench (Zheng et al. 2023): MT-Bench is specifically designed to evaluate LLMs’ capability as evaluators. The benchmark incorporates pairwise comparison data for 80 questions. Each pairwise data receives multiple human annotations, resulting in a total of 3k votes. To ensure high-quality evaluation samples, we select only those cases where all human annotators reach complete consensus, ultimately obtaining 684 evaluation samples.
  • WildVision (Lu et al. 2024): WildVision comprises 11k human-annotated preference relations among LMM response pairs. Each relation includes a question-image pair and two responses generated by different models, with a human-annotated preference (including ties). For testing, we randomly select 2k response pairs from the dataset using the same testing protocol as LLaVA-Critic.
  • MLLM-as-a-Judge (Chen et al. 2024): MLLM-as-a-Judge establishes a novel framework for assessing how closely model evaluations align with human judgments. The benchmark aggregates 17k multimodal evaluation instances (image-instruction-response triplets) spanning 14 diverse benchmarks and incorporating outputs from 6 different LMMs. Through systematic pairwise comparisons conducted by human evaluators, the dataset provides 5,719 carefully annotated judgment cases for analysis.
  • VLRewardBench (Li et al. 2024c): VLRewardBench is a comprehensive benchmark spanning general multimodal queries, visual hallucination detection, and complex reasoning tasks. Through an AI-assisted annotation pipeline combining sample selection with human verification, it curates 1,250 high-quality examples specifically designed to probe model limitations.
  • GenAI-Bench (Li et al. 2024a): GenAI-Bench is a comprehensive benchmark with 1.6k compositional prompts to evaluate text-to-visual generation, surpassing the size and difficulty of existing benchmarks. Additionally, it provides over 15k human ratings for multiple aspects to support research on vision-language alignment metrics.
  • Winoground (Thrush et al. 2022): Winoground is a benchmark for evaluating vision-language models’ ability to perform visio-linguistic compositional reasoning—the capacity to understand how the meaning of an image changes when the same words are rearranged in a caption. It was hand-curated by expert annotators and labeled with fine-grained tags, containing 400 samples.
  • Pickapic (Kirstain et al. 2023): Pickapic is a large dataset of text-to-image prompts and real users’ preferences over generated images. Authors create a web app that enables text-to-image users to generate images and specify their preferences, which constructs 500 test samples.
Aspects Used in the public meta-benchmark
Since UFEval requires aspects for evaluation, we assign one to each benchmark. Specifically, we select suitable aspects from the TAs and UAs, and for unseen tasks, we define appropriate aspects based on task properties. Tables 24 to 27 show the aspects used by UFEval, with blue indicating unseen aspects. GPT-4o, Claude-3.5, Qwen2VL-7B, and Qwen2VL-72B also follow the same aspect assignments.
@@ Appendix C Appendix C: Experiment Detail
Table 6: The detailed statistical information of the sub-tasks is not shown in Figure 2 of the main text.
@@
Table 7: Different LMMs used for generating pairwise responses for sub-tasks under the “Generated” column in Figure 2 of the main text.

### v4

@@ 4 Experiments
To assess UFEval’s generalizability from inter-aspect correlations, we first conduct evaluations on previously unseen tasks and aspects, termed Out-of-Domain Evaluation. Next, we evaluate UFEval as MLLM-as-a-Judge using public benchmarks. We further explore the effectiveness of Multi-aspect and Multi-task Assessment Learning for evaluators. Lastly, we validate UFEval’s application in generating preference data for DPO-based alignment. In the following sections, we present our experimental setup and results (In-Domain Evaluation and other experimental results are in Appendix D).
@@
Figure 3: Comparison of baselines on the FRA-OOD and FRA-OOD-H. Black areas show unseen TAs (for aspect generalization), while the rest show seen UAs for unseen tasks (for task generalization).
Table 2: The results correspond to task and aspect generalization evaluation. Average accuracy serves as the evaluation metric. Bold and underline indicate the first and second best results, respectively.
Method Task Generalization Evaluation Aspect Generalization Evaluation
@@
4.1 Experimental Setup
4.1.1 Benchmarks.
(1) Out-of-Domain Evaluation: we use FRA-OOD and FRA-OOD-H to validate. (2) Evaluation as MLLM-as-a-Judge: we select public benchmarks across three tasks: For NLG, we use MT-Bench (Zheng et al., 2023), SummEval (Fabbri et al., 2021), and MANS (Guan et al., 2021). Since SummEval and MANS only provide scores, we generate pairwise samples for evaluation through sample pairing. For IU, we use WildVision (Lu et al., 2024), MLLM-as-a-Judge (Chen et al., 2024), and VLRewardBench (Li et al., 2024e). For IG, GenAI-Bench (Li et al., 2024a), Winoground (Thrush et al., 2022), and Pick-a-Pic (Kirstain et al., 2023) are selected. To ensure fair comparison, we carefully check that the training set contains no overlapping samples with the benchmarks. (3) Multi-Aspect and Multi-Task Assessment Learning: We use FRA-ID to investigate the multi-aspect learning and benchmarks above to verify multi-task synergy. (4) Preference Alignment Comparison: We leverage UFEval for image generation and understanding model alignment. For image generation, we generate images using captions from the HPDv2 (Wu et al., 2023). For image understanding, we use MMHal (Sun et al., 2023), LLaVABen.Wild (Li et al., 2024b), and LLaVABen (Liu et al., 2023a). Detailed descriptions and usage of each benchmark are in Appendix C.2.
Table 3: Evaluation as MLLM-as-a-Judge for NLG, using three benchmarks. “tau” indicates that accuracy is calculated with ties, and “diff” excludes tied pairs when calculating accuracy.
@@ 4.1 Experimental Setup
Table 4: Evaluation as MLLM-as-a-Judge for IU. We evaluate baselines across three benchmarks.
@@
4.1.2 Baselines.
We experiment with representative models from two categories: (1) Prompting Models: GPT-4o, Claude-3.5, Qwen2-VL-72B-Instruct (Qwen2VL-72B), and Qwen2-VL-7B-Instruct (Qwen2VL-7B), implemented using the same prompts as UFEval. (2) Fine-tuned Evaluators: Themis-8B (Themis) (Hu et al., 2024b) , Auto-J (Li et al., 2023b), Prometheus 2 (Kim et al., 2024a), LLaVA-Critic-7B (LLaVA-Critic) (Xiong et al., 2024), ImageReward (Xu et al., 2023a), and VisionReward (Xu et al., 2024). Notably, ImageReward and VisionReward use different evaluation paradigms from ours. Auto-J cannot assess specific aspects, and Prometheus 2 requires reference answers. Therefore, we use these methods only as baselines in the MLLM-as-a-Judge evaluation.
4.2 Out-of-Domain Evaluation
Tasks Generalization Evaluation. For task generalization, we employ the seen UAs and unseen tasks from UFEval during testing to control for test variables. The results are shown in the regions outside the black sectors in Figure 3  [^3] We use aspect abbreviations here; the full names can be found in Appendix A.. Detailed scores for each aspect are reported in Appendix D.4. UFEval demonstrates strong alignment with both GPT-4o and human annotators on all tasks. In the “Tasks Generalization Evaluation” column of Table 2, UFEval achieves overall average accuracies of 85% and 83% on FRA-OOD and FRA-OOD-H, respectively, and outperforms both Themis and LLaVA-Critic on all tasks. Supporting the effectiveness of aspect-level evaluation for cross-task generalization. We also provide several good and bad cases in Appendix H.
Aspects Generalization Evaluation. We use the unseen aspects of unseen tasks in FRA-OOD and FRA-OOD-H to evaluate UFEval’s generalization to aspects. To support the validity of aspect generalization evaluation, we provide ROUGE_L (Wang et al., 2022) tests on unseen aspects in Appendix B.3. The results are shown in the black sectors of Figure 3. In terms of overall coverage, UFEval outperforms both Themis and LLaVA-Critic, the state-of-the-art evaluators in their respective domains. In the right column of Table 2, UFEval continues to demonstrate strong performance on the NLG, IU, IG and ITIG, with high consistency with both GPT-4o and human annotators, achieving overall accuracies of 86.2% and 83.2%, respectively. These results show that even without exposure to unseen aspects during training, UFEval remains effective in evaluation. This is largely due to the naturally related nature of aspects, which enables the transfer of similar meanings and standards.
4.3 Evaluation as MLLM-as-a-Judge
Table 5: Evaluation as MLLM-as-a-Judge for IG. We evaluate baselines across three benchmarks.
Method GenAI-Bench Winoground Pick-a-Pic
@@ 4.3 Evaluation as MLLM-as-a-Judge
Figure 4: The results of multi-aspect assessment learning. We train models using different training set combinations and test them on three visual tasks (IU, IG, and ITIG) in FRA-ID to validate the effectiveness of multi-aspect learning.
For NLG Evaluations. The experimental results, shown in Table 3, demonstrate the effectiveness of UFEval in NLG tasks. Specifically, in SummEval and MT-Bench, UFEval outperforms most baselines, with Claude being the sole exception. In the MANA benchmark, UFEval achieves the best performance, with an accuracy of 69.3%. These results highlight UFEval’s strong capability for text understanding evaluation compared to domain-specific models in this field.
For IU Evaluations. The results are presented in Table 4. Additionally, we assess model-level consistency on WildVision using Elo ratings and Kendall’s Tau ($\tau$) to compare model rankings (Xiong et al., 2024). The findings demonstrate that UFEval consistently outperforms LLaVA-Critic across all benchmarks. LLaVA-Critic focuses on respective datasets training, whereas UFEval uses joint learning across multiple tasks, which leads to relatively superior performance. For more comprehensive MLLM-as-a-Judge results, please refer to Table 48.
For IG Evaluations. The experimental results are shown in Table 5. Our evaluator outperforms ImageReward in all benchmarks, achieving accuracies of 65.5%, 65.7%, and 57.3%, respectively. Even compared to the state-of-the-art image evaluator VisionReward, the performance gap is minimal—only 0.9% lower on GenAI-Bench and 0.3% lower on Winogrounded. These results demonstrate the promising capability of UFEval in the image generation evaluation task.
4.4 Multi-aspect and Multi-task Assessment Learning
@@
Table 6: Image understanding DPO comparison. We compare our UFEval with LLaVA-Critic for DPO based on LLaVA-Next-7B.
Method MMHal LLaVABen.Wild LLaVABen
LLaVA-Next-7B 2.05 52.9 28.3
w/ LLaVA-Critic 2.24 59.0 30.3
w/ UFEval 2.41 61.4 32.3
Table 7: Image generation DPO comparison. We evaluate DPO performance using UFEval-generated data versus the Pick-a-Pic dataset.
Method HPSv2 ImageReward VisionReward
SDXL 28.1 0.80 3.00
w/ Pick-a-Pic 28.7 0.84 3.20
w/ UFEval 29.9 0.90 3.27
4.5 Preference Alignment Comparison
To validate the effectiveness of UFEval to generate preference data across both IU and IG tasks, we employ it to construct training data for DPO-based model alignment. The application principle of DPO for model alignment in both IU and IG tasks is comprehensively described in Appendix F.
For IU Tasks. Building upon UFEval, we use DPO to improve the image understanding capabilities of LLaVA-Next-7B (Li et al., 2024d). For fair comparison, both UFEval and the LLaVA-Critic ues identical image-question pairs sourced from RLHF-V (Yu et al., 2024) and LLaVA-RLHF (Sun et al., 2023) to construct preference data. Ultimately, we construct 15k preference samples. We then train LLaVA-Next-7B on 8 A100 GPUs using a batch size of 2, gradient accumulation steps of 2, a learning rate of $5\times 10^{-7}$, and set $\beta_{u}$ to 0.1. The results are shown in Table 7, UFEval consistently surpasses LLaVA-Critic across all evaluated benchmarks. Notably, it achieves a 2.4% improvement on LLaVABench.Wild, underscoring its enhanced effectiveness in visual understanding tasks.
For IG Tasks. Based on UFEval, we apply DPO to SDXL-Turbo (Podell et al., 2023), a conditional diffusion model, to better align its outputs with human preferences without explicit reward modeling. We extract prompt and two corresponding images from Pick-a-Pic and use UFEval to construct training preference data. In total, we construct 14k preference samples for DPO image generation. We then train SDXL-Turbo on 8 A100 GPUs using $\beta_{g}=5000$ with a batch size of 32 for three epochs. HPSv2 (Wu et al., 2023), ImageReward, and VisionReward are used for quality assessment. The results are shown in Table 7, training on the constructed data using UFEval achieves better performance compared to directly training on the original dataset. The qualitative comparison results are shown in Appendix I. This demonstrates the effectiveness of our approach in refining preference data for improved model alignment in image generation tasks.
5 Conclusion
This paper proposes UFEval, the first unified fine-grained evaluator with task and aspect generalization. Specifically, we start by building a comprehensive aspect tree that leads to the creation of FRABench, a large-scale, multi-modal, and aspect-level evaluation dataset. We then fine-tune an MLLM on FRABench to develop UFEval. Our experimental results demonstrate that joint learning across diverse visual tasks and aspects yields significant mutual benefits and generalization capabilities. We also leverage UFEval to automatically construct high-quality preference pair datasets for DPO training to align models’ outputs. These results validate the effectiveness of the FRABench as a valuable resource for training unified evaluators. In future work, we plan to incorporate video understanding and generation tasks into our evaluation system and add their corresponding aspects to the aspect tree.
Appendix A Aspect Sources and Definitions
We provide the sources and definitions of aspects under two categories: UAs and TAs, as shown in Tables 12 to 22. For UAs, the “Target” column indicates the applicable output modality: “T’ for Text, “I” for Image, and “ITI” for Interleaved Text-with-Image. For TAs, the “Sub-Task” column specifies the corresponding task type. Additionally, underscores are used to benchmark extended TAs for ITIG. Abbreviations for all aspects are provided in the table.
Appendix B Analysis of the FRABench
@@
B.1 More Detailed Information about the FRABench
@@
We provide the information about sub-tasks for NLG and IU in Table 8, as this information is not shown in Figure 2 of the main text. For the tasks in the “Generated” column in Figure 2 of the main text, due to the absence of pairwise data, we use four MLLMs with varying performance levels to generate pairwise responses for each query, as illustrated in Table 9.
Table 8: The detailed statistical information of the sub-tasks, which is not shown in Figure 2.
@@
Table 9: Different MLLMs used for generating pairwise responses for sub-tasks under the “Generated” column in Figure 2 of the main text.
@@
B.2 Position Bias Analysis
B.3 Aspect Diversity Analysis
Following prior work (Wang et al., 2022; Honovich et al., 2022), we analyze the ROUGE-L distribution between aspects in the training set and the FRA-OOD, respectively. Specifically, we sample pairs of aspects from the training set and compute their ROUGE-L scores. The overall distribution is shown in Figure 9, which indicates that the selected aspects are distinct from one another, confirming the inclusion of diverse and novel aspects in the training set. Additionally, we measure the ROUGE-L scores between aspects by sampling one aspect from the training set and another from the FRA-OOD. As shown in Figure 10, aspects in the training set and those in the FRA-OOD do not overlap, demonstrating the validity of our out-of-domain evaluations.
B.4 Aspect Selection Across Different Sub-tasks
We provide detailed information on the aspects selected for the training set, FRA-ID, FRAUAs-OOD, and FRA-OOD, as presented in Table 11. Additionally, we include statistics regarding the number of evaluation labels used for training and testing in each sub-task. Blue-marked aspects are human-annotated, while the rest are annotated by GPT-4o. The human-annotated data in IG is sourced from the ImageRewardDB dataset (Xu et al., 2023a).